OpenStudy (anonymous):

Compute the monthly payments for an add-on interest loan of $850, with an annual interest rate of 15 percent and a term of 4 years. Round to the nearest cent as needed.

OpenStudy (anonymous):

payments = A/n where, A = P + I,= 850 +127.50 = 977.50 I = Pr = 850x.15 = 127.5 n= months x years= (in this case 12x4=48) so, A/n =977.50/48 = 20.36